# Mario Shannon

### v1

@@ 2.5. Construction of the paths
To derive (4) from (3) we have to compare the periods of a $\phi_{t}^{A}$-periodic orbit $\alpha$ and the corresponding $\psi_{t}$-periodic orbit $\beta=\mathcal{P}(\alpha)$. To do this, let $T$ be a global transverse section of $\phi_{t}^{A}:M_{A}\to M_{A}$. The surgery $(\phi_{t}^{A},M_{A})\mapsto(\psi_{t},N)$ transforms the global transverse section $T$ in a Birkhoff section for $(\psi_{t},N)$ that we denote by $\iota:(S,\partial S)\to(N,\{\delta\})$. It has $p(S,\delta)=\textsl{per}_{\mathbb{Z}}(\gamma)$ boundary components (cf. definition 18) attached along $\delta$ with multiplicity $m$. It is direct to check that every $\phi_{t}^{A}$-periodic orbit $\alpha\neq\gamma$ is sent to a $\psi_{t}$-periodic orbit $\beta$ transverse to the Birkhoff section $S$ satisfying $\textsl{per}_{\mathbb{Z}}(\alpha)=[\beta]\cdot[S]$, where this last expression indicates pairing of relative homology classes $H_{*}(N;\delta)$. We have that:
@@
  • Since the $\phi_{t}^{A}$-orbits are transverse to $T$ there exists $\tau_{1}>0$ such that every $\phi_{t}^{A}$-orbit segment of length at most $\tau_{1}$ intersects at most once the surface $T$. Hence, every periodic orbit $\alpha$ with period $\textsl{per}(\alpha)=t>\tau_{1}$ has a number of intersections with $T$ bounded by $t/\tau_{1}$. It follows that:
@@
    $\textsl{per}(\alpha)\geq\kappa_{1}\cdot\textsl{per}_{\mathbb{Z}}(\alpha),\ \text{for some constant}\ 0<\kappa_{1}<\tau_{1}.$
@@
  • There exists $\tau_{2}>0$ such that every $\psi_{t}$-orbit segment of length bigger that $\tau_{2}$ has at least one intersection with the Birkhoff section $S$. Hence, every $\psi_{t}$-periodic orbit $\beta$ with period $\textsl{per}(\beta)=t>\tau_{2}$ and transverse to the interior of $S$ must intersect the surface in at least $t/\tau_{2}$ points. It follows that:
@@
    $[\beta]\cdot[S]\geq\kappa_{2}\cdot\textsl{per}(\beta),\ \text{for some constant}\ 0<\kappa_{2}<1/\tau_{2}.$
@@
From the items above we see that there exists $\kappa_{3}>0$ such that $\textsl{per}(\alpha)\geq\kappa_{3}\cdot\textsl{per}(\beta)$, for every $\phi_{t}^{A}$-periodic orbit $\alpha$ with period $\textsl{per}(\alpha)>\tau_{1},\tau_{2}$. This implies that
[…]
and hence, for some constants $C_{1}>C_{0}\sqrt{1/\kappa_{3}}$ and $t_{1}>t_{0}$ we have that
@@
|$|\{\alpha\in\mathcal{Q}(\gamma,m):\textsl{per}(\alpha)\leq t\}|\leq C_{0}\sqrt{t/\kappa_{3}}\log(t/\kappa_{3})\leq C_{1}\sqrt{t}\log(t),\ \forall\ t>t_{1}.$
@@ Acknowledgements
We want to thank S. Fenley for useful comments and discussions around the present result.

### v2

@@ 2.5. Construction of the paths
transforms the global transverse section $T$ into a Birkhoff section for $(\psi,N)$ that we denote by $\iota:(S,\partial S)\to(N,\{\delta\})$. It has $p(S,\delta)=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\gamma)$ boundary components (cf. Definition 20) attached along $\delta$ with multiplicity $m$. It is direct to check that every $\phi^{A}$-periodic orbit $\alpha\neq\gamma$ is sent to a $\psi$-periodic orbit $\beta$ transverse to the Birkhoff section $S$ satisfying $\beta]\cdot[S]=\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha)$, where this last expression indicates pairing of relative homology classes in $H_{*}(N;\delta)$. Furthermore, since the manifolds supporting the flows $\phi^{A}$ and $\psi$ are compact, there exists constants $0<a_{1}<a_{2}$ and $0<b_{1}<b_{2}$ such that
$a_{1}\cdot\operatorname{\mathrm{per}}(\alpha)\leq\operatorname{\mathrm{length}}(\alpha)\leq a_{2}\cdot\operatorname{\mathrm{per}}(\alpha),\ \text{for every}\ \alpha\in\textrm{Per}(\phi^{A},M_{A})$
$b_{1}\cdot\operatorname{\mathrm{per}}(\beta)\leq\operatorname{\mathrm{length}}(\beta)\leq b_{2}\cdot\operatorname{\mathrm{per}}(\beta),\ \text{for every}\ \beta\in\textrm{Per}(\psi,N).$
We have that:
@@
  • Since the $\phi^{A}$-orbits are transverse to $T$ there exists $l_{1}>0$ such that every $\phi^{A}$-orbit segment of length at most $l_{1}$ intersects at most once the transverse section $T$. Hence, every periodic orbit $\alpha$ with length bigger than $l_{1}$ has a number of intersections with $T$ bounded from above by $(a_{2}/l_{1})\cdot\operatorname{\mathrm{per}}(\alpha)$. It follows that:
@@
    $\operatorname{\mathrm{per}}(\alpha)\geq\kappa_{1}\cdot\operatorname{\mathrm{per}}_{\mathbb{Z}}(\alpha),\ \text{for some constant}\ \kappa_{1}>0.$
@@
  • There exists $l_{2}>0$ such that every $\psi$-orbit segment of length bigger that $\tau_{2}$ has at least one intersection with the Birkhoff section $S$. Hence, every $\psi$-periodic orbit $\beta$ with length bigger that $l_{2}$ and transverse to the interior of $S$ must intersect this surface in at least $(b_{1}/l_{2})\cdot\operatorname{\mathrm{per}}(\beta)$ points. It follows that:
@@
    $\beta]\cdot[S]\geq\kappa_{2}\cdot\operatorname{\mathrm{per}}(\beta),\ \text{for some constant}\ \kappa_{2}>0.$
@@
In addition, since every Anosov flow has at most finitely many periodic orbits of length (or period) lower than any given constant, there exists $t_{*}>0$ such that $\operatorname{\mathrm{length}}(\alpha)>l_{1}$ and $\operatorname{\mathrm{length}}(\beta)>l_{2}$ whenever $\operatorname{\mathrm{per}}(\alpha)>t_{*}$. Hence, from the items above we see that there exists $\kappa_{3}>0$ such that $\operatorname{\mathrm{per}}(\alpha)\geq\kappa_{3}\cdot\operatorname{\mathrm{per}}(\beta)$, whenever $\operatorname{\mathrm{per}}(\alpha)>t_{*}$. This implies
$\{\alpha\in\mathcal{Q}(\gamma,m):t_{*}<\operatorname{\mathrm{per}}(\alpha)\leq t\}\subset\{\Pi^{-1}(\beta):\beta\in\mathcal{FH}(\delta)\text{ and }\operatorname{\mathrm{per}}(\beta)\leq t/\kappa_{3}\},\ \forall\ t>t_{*}.$
Finally, by setting $C_{*}=|\{\alpha\in\mathcal{Q}(\gamma,m):\operatorname{\mathrm{per}}(\alpha)\leq t_{*}\}|$, there exists constants $C_{1}>C_{0}\sqrt{1/\kappa_{3}}$ and $t_{1}>\max\{t_{0},t_{*}\}$ such that
@@
$|\{\alpha\in\mathcal{Q}(\gamma,m):\operatorname{\mathrm{per}}(\alpha)\leq t\}|\leq C_{*}+C_{0}\sqrt{t/\kappa_{3}}\log(t/\kappa_{3})\leq C_{1}\sqrt{t}\log(t),\ \forall\ t>t_{1}.\qed$
@@ Acknowledgements
We want to thank S. Fenley for useful comments and discussions on the present result.